An earth satellite has a mass M and angular momentum L.Its areal velocity is given by........

Grade:Select Grade

1 Answers

Aziz Alam IIT Roorkee
9 years ago
L=mvr

Area covered by satellite at any given point is equal to,

A=\pi R^{2}\theta /2\pi =R^{2}\theta /2
dA/dt=\omega R^{2}/2=vR/2
vR/2=L/2m